### v1

@@ Abstract
Research in various fields is currently experiencing challenges regarding the reproducibility of results. This problem is also prevalent in machine learning (ML) research. The issue arises primarily due to unpublished data and/or source code and the sensitivity of ML training conditions. Although different solutions have been proposed to address this issue, such as using ML platforms, the level of reproducibility in ML-driven research remains unsatisfactory. Therefore, in this article, we discuss the reproducibility of ML-driven research with three main aims: (i) identify the barriers to reproducibility when applying ML in research as well as categorize the barriers to different types of reproducibility (description, code, data, and experiment reproducibility), (ii) identify potential drivers such as tools, practices, and interventions that support ML reproducibility as well as distinguish between technology-driven drivers, procedural drivers, and drivers related to awareness and education, and (iii) map the drivers to the barriers. With this work, we hope to provide insights and contribute to the decision-making process regarding the adoption of different solutions to support ML reproducibility.
@@ 1 Introduction
As in many scientific fields (e.g., [8, 66]), research in artificial intelligence (AI) in general, and machine learning (ML) in particular, faces crucial doubts over the reproducibility of research [35]. This has raised concerns about the reliability and validity of many scientific findings, risking diminishing confidence in the overall body of scientific knowledge. Misleading or irreproducible results can lead to wasted resources, hinder scientific progress, reduce trust in science, and impact decision-making in various fields [49, 28]. The “reproducibility crisis” has been a topic of discussion since the mid-2010s, particularly in the medical and behavioral science fields [8, 53]. Recently, this issue has also been associated with research involving ML methods [35]. Apart from the common challenges faced by many disciplines, such as limited sharing of data and source code, the use of ML introduces unique challenges for reproducibility, including sensitivity to ML training conditions, sources of randomness [63], and the increasing use of AutoML tools [41].
While the literature on these issues is expanding [45, 29, 32, 39, 25, 4, 68, 26], there is still no comprehensive overview of the associated barriers and drivers. For example, in [26], the authors identify and categorize sources of irreproducibility in ML and how these sources affect conclusions drawn from ML experiments. However, this study does not investigate drivers to address these sources of irreproducibility. Thus, our paper provides a contextual categorization of the barriers and drivers to the four types of ML reproducibility proposed by [25], with specific reference to research in both Computer Science and Biomedical fields. We also propose a Drivers-Barriers-Matrix to summarize and visualize the results of the discussion. Such an analysis stands to clarify the current state regarding ML reproducibility, provide guidance for those new to the subject, and highlight potential gaps, solutions, and avenues for future exploration.
@@ 2 Defining Reproducibility
Within AI research, widely accepted definitions that build further on these key distinctions have been proposed by  [29, 25]. We follow and build upon these latter definitions, and hence here outline them in some length.
@@
This categorization aims to overcome the problem of ambiguity when making specific claims about the reproducibility of an experiment. Often in literature, authors write about reproducing the same “results” of an experiment. It is not apparent, however, in which cases they mean to achieve the same computational outcome, i.e., outputs of the algorithms, or whether they mean to reach the same analysis or interpretation. Therefore, achieving interpretation reproducibility is a more general and often less stringent requirement than achieving outcome reproducibility. This categorization is not specific to ML but generally applicable to any research field that conducts computational experiments.
In addition, relating to point (i) of the Goodman et al. schema (“methods” reproducibility), [25] specifies degrees to which methods can be made transparent through description or sharing. These four types are defined as R1 Description, R2 Code, R3 Data, and R4 Experiment. The lower the level of reproducibility, the less shared information there is, making the experiment more difficult to reproduce. Giving an example: In general, all published research experiments are accompanied by a textual description of the experiment. If this textual description is the only information shared by the authors, the research is categorized as R1 Description, which, according to this scheme, is the lowest type of reproducibility. In contrast, if all three building blocks, i.e., text, code, and data, are shared, the experiment can be categorized as R4 Experiment, which is the highest type of reproducibility. Furthermore, the distinction between R2 Code and R3 Data is defined by the textual description of the original experiment that is accompanied by either code or data, respectively.
Also, the different types of ML reproducibility exhibit an interplay between generalizability and reproducibility. R1 Description leads to the highest generalizability but to the weakest reproducibility, while R4 Experiment leads to the highest reproducibility and lowest generalizability.
This relationship is illustrated in Figure 1 [25]. Furthermore, Figure 1 visualizes the building blocks associated with the different types of reproducibility. In what follows, unless stated otherwise, we use the definitions proposed by Gundersen and colleagues.
@@ 3.1 R1 Description
    Often results are selectively reported, e.g., researchers may only provide results for the best performance out of many test runs, instead of properly assessing and reporting the performance over multiple runs. For example, in research on Natural Language Processing (NLP), often, reproductions turn out to perform worse than stated in publications [10].
Generally, it is important for studies to use a robust methodology and provide detailed reports, for other researchers to verify results and understand how analyses were conducted. While ML models have proven highly effective in Biomedical fields, studies often fall short in providing comprehensive and high-quality reporting. For example, in studies of predicting cardiometabolic risk from dietary patterns [54] or supporting the clinical management of diabetes [38], ML models were observed to be very promising. While this further enhances the promise of applying ML models for various clinical prediction tasks, there is a clear need for thorough reporting and validation of these models to allow for their integration into routine clinical care [38]. This is also true for the application of ML models for cancer imaging, where ML models often surpass radiologists in performance, but publications on these models lack the details needed to reproduce the results [59].
@@
Another flaw, which has been observed in ML-based research and negatively affects reproducibility, is “spin”. It refers to the misuse of language to “intentionally or unintentionally affect the interpretation of study findings”. Accordingly, it is also understood as an inconsistency between the study results and the conclusions, in the sense that results are over-generalized or the claimed conclusions are not supported by the scientific method. This has been shown to impact both the interpretations and decision-making by readers [5]. In ML-based Biomedical research, the most common practice of spin includes recommending models for various applications without providing external validation in the same study. More concretely, the recommendation to use a model either in a clinical setting or for a different population is only validated in approximately 15% of the cases. Other observed instances of spin are invalid comparisons of results to previous developments and the use of leading words and strong statements to make the results sound more prolific [5]. The use of spin in numerous publications might be partly attributed to the reward system in academia, which often values spin practices in research, as more prolific results are more likely to lead to accepted publications. Publication bias in general is a big problem in a variety of different regards. It seems that research from high-profile institutions is published more easily and subject to less scrutiny in review and there also seems to be a tendency to prefer more complex solutions over simpler ones, which solve the same problem [72].
@@ 3.2 R2 Code
Published ML research is often not accompanied by available data and code. Only one-third of researchers share data, and even fewer share their source code [36]. This can be for many reasons, such as the increasing pressure on researchers to publish quickly, which in turn does not allow them to polish the code and decreases the willingness to release the code, or simply due to private source code that builds on unpublished code. Sharing source code correctly requires publishing seven pieces of information: hypothesis, prediction method, source code, hardware specifications, software dependencies, experiment setup, and experiment source code. Only if the source code is accompanied by all these artifacts, it is specified sufficiently to allow reproducing an experiment [28]. Unfortunately, current research often does not meet this requirement, leading to reproducibility issues due to different software versions, hyperparameter settings, or hardware differences  [33, 10]. For example, in Computer Science, the lack of shared code and data significantly contributes to a lack of reproducibility [18]. Even when code and data are shared, they are often poorly documented, or the code is only provided as a skeleton instead of a fully executable code, which does not ensure reproducibility. Thus, shared code should also include the code for data preprocessing, hyperparameter tuning, and managing random seeds, as well as code for comparisons against baseline models.
@@ 3.3 R3 Data
The main barrier concerning R3 Data is that data is often not shared. This could be due to privacy concerns or a lack of incentives and motivation. Moreover, many benchmark and training datasets encounter challenges related to copyright, licensing, and longevity. These datasets may also raise ethical concerns, such as the unintentional inclusion of privacy-sensitive or harmful content, making it difficult to share the data for ML model training [55]. Similar to the sharing of source code, sharing only the datasets is insufficient. For proper use, it is also important to share specific splits, i.e., the training dataset, validation dataset, and test dataset [28]. Furthermore, data sharing needs to be accompanied by details about the provenance and preprocessing of data. This is crucial to investigate methodological errors such as data leakage and bias, which are described in the following.
@@
In practice, methodological issues such as data leakage often hinder the reproducibility of ML-based research [39]. This is due to the growing number of non-experts employing machine learning across different research fields [23], which is fueled by the ease of application of ML libraries and no-code off-the-shelf AI tools. In essence, data leakage happens when data on which the ML model should not be trained leaks into the training process. Data leakage can be categorized into 3 subcategories [39]:
@@
    Use of non-legitimate data. For example, when the use of antihypertensive drugs is used to predict hypertension. This data is non-legitimate since it would not be available in a real-world scenario and cannot realistically be used to predict hypertension for a new patient.
@@
    Test set is not drawn from the distribution of scientific interest. There are three possible variants: (1) temporal leakage, which is problematic for ML models that attempt to predict future outcomes, i.e., when some training samples have a later timestamp than samples available in the test set. (2) the training and test data are not independent of each other, e.g., there should not be samples in the training and test data which are drawn from the same person. (3) the test set is not chosen selectively; for instance, if the model is solely evaluated on data where it performs well.
@@
Bias in ML refers to the error introduced by approximating a real-world problem, which may be complex, by a simplified model, often leading to systematic deviations from the ground truth. Furthermore, this bias can arise when the model contains imbalances or reflects existing societal biases. ML models, which are subject to bias, do not generalize well and are therefore potentially problematic for ML reproducibility. There are eight bias variants that can arise during the data handling phase of ML development [65]: (1) selection bias - using data not being representative of the target group, (2) exclusion bias - excluding particular data samples based on the belief that they are unimportant, (3) measurement bias - favoring certain measurement results, (4) recall bias - labeling similar data samples differently, (5) survey bias - introducing data issues stemming from data collection surveys, (6) confirmation bias - favoring information, which confirms previous beliefs, (7) prejudice bias - including human-related prejudices in training data, and (8) algorithmic bias - replicating or amplifying biases by algorithms themselves. Biases, such as selection bias, often lead to the issue of validity shrinkage in Biomedical science research [37]. For example, in obesity and nutritional research, ML is used to predict obesity, heart rate, or the risk of heart attack, based on data of an individual. Here, validity shrinkage refers to the issue that a predictive model trained on a subset of data will most probably not perform well on new samples. The difference between predictive performance on known data and new data, however, is most often not accounted for in nutritional science and therefore also leads to performance claims that cannot be reproduced [37].
@@ 3.4 R4 Experiment
Inherent nondeterminism in ML models means that results can vary between test runs, even with identical code, data, and hyperparameters. This variation occurs from unanticipated influences during training, such as random parameter initialization, stochastic optimization, and random data subsampling (e.g., in k-fold cross-validation) [63, 43]. Neural networks are especially known for their inherent nondeterminism, leading to varied computational outcomes in multiple reruns due to numerous sources of randomness during training [1]. In some cases, inherent non-determinism can cause such large variations that reruns not only yield slightly different outcomes but also lead to significant fluctuations in the performance of an ML model [1]. This issue is exacerbated when other sources of variation, such as different hyperparameters, are introduced to the ML model. In such cases, the impact can be magnified, and it is often observed that minor changes in hyperparameters can result in significant performance loss [10]. For example, a review of the status quo of reproducibility in NLP research highlights some core issues with non-determinism that are exemplary for ML research [10]. Simply rerunning the original code during a reproduction can lead to different computational outcomes as the original reported ones. Reinforcement learning (RL), a subfield of ML, is particularly susceptible to these reproducibility issues, partially because of additional sources of non-determinism in the learning process, such as the RL environment or policy [50].
@@
Various studies have demonstrated that hardware differences, such as different GPUs or CPUs, and compiler settings can lead to different computational outcomes [33]. Additionally, a comparison between the same ML algorithm with fixed random seeds executed using PyTorch [^1] https://pytorch.org/ and TensorFlow [^2] https://www.tensorflow.org/ resulted in different performances [58]. A comparison of the results of an ML experiment between different ML platforms also shows that out-of-the-box reproducibility is not provided. In general, it is questionable to what extent reproducibility can be ensured out-of-the-box [29]. Another important factor is the use of GPUs, which can increase randomness compared to the use of CPUs. This is due to parallel optimization and the use of optimizers in deep learning frameworks, such as PyTorch and TensorFlow. As a result, some researchers have resorted to solely using CPUs for computation. However, this comes at the expense of time-efficiency [3].
@@
Another barrier to ML reproducibility is the limited access to computational resources, which most recently has become evident in the case of transformer-based large language models (LLMs) [9, 79]. These transformer architectures need a vast amount of data and computational resources, to which most researchers have limited access. Estimates have calculated the costs to reproduce one model to be around \$1 million to \$3.2 million, which largely increases the difficulty of reproducing results [9]. Another study found that the needed computational resources are one of the most significant factors impacting reproducibility [62]. Especially ML models, which require computational clusters for training, are notably hard to reproduce.

### v2

@@ Abstract
Research in various fields is currently experiencing challenges regarding the reproducibility of results. This problem is also prevalent in machine learning (ML) research. The issue arises, for example, due to unpublished data and/or source code and the sensitivity of ML training conditions. Although different solutions have been proposed to address this issue, such as using ML platforms, the level of reproducibility in ML-driven research remains unsatisfactory. Therefore, in this article, we discuss the reproducibility of ML-driven research with three main aims: (i) identifying the barriers to reproducibility when applying ML in research as well as categorize the barriers to different types of reproducibility (description, code, data, and experiment reproducibility), (ii) discussing potential drivers such as tools, practices, and interventions that support ML reproducibility, as well as distinguish between technology-driven drivers, procedural drivers, and drivers related to awareness and education, and (iii) mapping the drivers to the barriers. With this work, we hope to provide insights and to contribute to the decision-making process regarding the adoption of different solutions to support ML reproducibility.
@@ 1 Introduction
As in many scientific fields [9, 68], e.g., medicine and behavior science, research in artificial intelligence (AI) in general, and machine learning (ML) in particular, faces crucial doubts over the reproducibility of research [36]. This has raised concerns about the reliability and validity of many scientific findings, risking diminishing confidence in the overall body of scientific knowledge. Misleading or irreproducible results can lead to wasted resources, hinder scientific progress, reduce trust in science, and impact decision-making in various fields [51, 29]. The “reproducibility crisis” has been a topic of discussion since the mid-2010s, particularly in the medical and behavioral science fields [9, 55]. Recently, this issue has also been associated with research involving ML methods [36]. Apart from the common challenges faced by many disciplines, such as limited sharing of data and source code, the use of ML introduces unique challenges for reproducibility, including sensitivity to ML training conditions, sources of randomness [65], and the increasing use of AutoML tools [42].
While the literature on these issues is expanding [46, 30, 33, 40, 26, 4, 70, 27], there is still no comprehensive overview of the associated barriers and drivers. For example, in [27], the authors identify and categorize sources of irreproducibility in ML and how these sources affect conclusions drawn from ML experiments. However, this study does not investigate drivers to address these sources of irreproducibility. Thus, our paper provides a contextual categorization of the barriers and drivers to the four types of ML reproducibility proposed by (description, code, data, and experiment) [26], with specific reference to research in both Computer Science and Biomedical fields. We also propose a Drivers-Barriers-Matrix to summarize and visualize the results of the discussion. Such an analysis stands to clarify the current state regarding ML reproducibility, provide guidance for those new to the subject, and highlight potential gaps, solutions, and avenues for future exploration.
@@ 2 Defining Reproducibility
Within ML research, widely accepted definitions that build further on these key distinctions have been proposed by  [30, 26]. We follow and build upon these latter definitions, and hence here outline them in some length.
@@
This categorization aims to overcome the problem of ambiguity when making specific claims about the reproducibility of an experiment. Often in literature, authors write about reproducing the same “results” of an experiment. It is not apparent, however, in which cases they mean to achieve the same computational outcome, i.e., outputs of the algorithms, or whether they mean to reach the same analysis or interpretation. Therefore, achieving interpretation reproducibility is a more general and often less stringent requirement than achieving outcome reproducibility. This categorization is not specific to ML, but generally applicable to any research field that conducts computational experiments.
In addition, relating to point (i) of the Goodman et al. schema (“methods” reproducibility), [26] specifies reproducibility types to which methods can be made transparent through description or sharing. These four types are defined as R1 Description, R2 Code, R3 Data, and R4 Experiment. The lower the level of reproducibility, the less shared information there is, making the experiment more difficult to reproduce. Giving an example: In general, all published research experiments are accompanied by a textual description of the experiment. If this textual description is the only information shared by the authors, the research is categorized as R1 Description, which, according to this scheme, is the lowest type of reproducibility. In contrast, if all three building blocks, i.e., text, code, and data, are shared, the experiment can be categorized as R4 Experiment, which is the highest type of reproducibility. Furthermore, the distinction between R2 Code and R3 Data is defined by the textual description of the original experiment that is accompanied by either code or data, respectively.
Also, the different types of ML reproducibility exhibit an interplay between generalizability and reproducibility. R1 Description leads to the highest generalizability, but to the weakest reproducibility, while R4 Experiment leads to the highest reproducibility and lowest generalizability.
This relationship is illustrated in Figure 1. Furthermore, Figure 1 visualizes the building blocks associated with the different types of reproducibility. In what follows, unless stated otherwise, we use the definitions proposed by Gundersen and colleagues [26].
@@ 3.1 R1 Description
    Often results are selectively reported, e.g., researchers may only provide results for the best performance out of many test runs, instead of properly assessing and reporting the performance over multiple runs. For example, in research on Natural Language Processing (NLP), often, reproductions turn out to perform worse than stated in the original publications [11].
Generally, it is important for studies to use a robust methodology and provide detailed reports, for other researchers to verify results and understand how analyses were conducted. While ML models have proven highly effective in Biomedical fields, studies often fall short in providing comprehensive and high-quality reporting. For example, in studies on predicting cardiometabolic risk from dietary patterns [56] or supporting the clinical management of diabetes [39], ML models were observed to be very promising. While this further enhances the promise of applying ML models for various clinical prediction tasks, there is a clear need for thorough reporting and validation of these models to allow for their integration into routine clinical care [39]. This is also true for the application of ML models for cancer imaging, where ML models often surpass radiologists in performance, but publications on these models lack the documentation details needed to reproduce the results [61].
@@
Another flaw, which has been observed in ML-based research and negatively affects reproducibility, is “spin”. It refers to the misuse of language to “intentionally or unintentionally affect the interpretation of study findings”. Accordingly, it is also understood as an inconsistency between the study results and the conclusions, in the sense that results are over-generalized or the claimed conclusions are not supported by the scientific method. This has been shown to impact both the interpretations and decision-making by readers [5]. In ML-based Biomedical research, the most common practice of spin includes recommending models for various applications without providing external validation in the same study. More concretely, the recommendation to use a model either in a clinical setting or for a different population is only validated in approximately 15% of the cases. Other observed instances of spin are invalid comparisons of results to previous studies, and the use of leading words and strong statements to make the results sound more prolific [5]. The use of spin in numerous publications might be partly attributed to the reward system in academia, which often values spin practices in research, as more prolific results are more likely to lead to accepted publications. Publication bias in general is a big problem in a variety of different regards. It seems that research from high-profile institutions is published more easily and subject to less scrutiny in review, and there also seems to be a tendency to prefer more complex solutions over simpler ones addressing the same problem [75].
@@ 3.2 R2 Code
Published ML research is often not accompanied by available data and code. Only one-third of researchers share data, and even fewer share their source code [37]. This can be for many reasons, such as the increasing pressure on researchers to publish quickly, which in turn does not allow them to polish the code and decreases the willingness to release the code, or simply due to privacy constraints with respect to sharing code. Sharing ML code correctly requires publishing seven pieces of information: hypothesis, prediction method, source code, hardware specifications, software dependencies, experiment setup, and experiment source code. Only if the ML code is accompanied by all these artifacts, it is specified sufficiently to allow reproducing an experiment [29]. Unfortunately, current research often does not meet this requirement, leading to reproducibility issues due to different software versions, hyperparameter settings, or hardware differences  [34, 11]. For example, in Computer Science, the lack of shared code and data significantly contributes to a lack of reproducibility [19]. Even when code and data are shared, they are often poorly documented, or the code is only provided as a skeleton instead of a fully executable code, which does not ensure reproducibility. Thus, shared code should also include the code and documentation for data preprocessing, hyperparameter tuning, and managing random seeds, as well as code for comparisons against baseline models.
@@ 3.3 R3 Data
The main barrier concerning R3 Data is that data is often not shared. This could be due to privacy concerns or a lack of incentives and motivation. Moreover, many benchmark and training datasets encounter challenges related to copyright, licensing, and longevity. These datasets may also raise ethical concerns, such as the unintentional inclusion of privacy-sensitive or harmful content, making it difficult to share the data for ML model training [57]. Similar to the sharing of source code, sharing only the datasets is insufficient. For proper use, it is also important to share specific splits, i.e., the training dataset, validation dataset, and test dataset [29]. Furthermore, data sharing needs to be accompanied by documentation specifying details about the provenance and preprocessing of data. This is crucial to investigate methodological errors such as data leakage and bias, which are described in the following.
@@
In practice, methodological issues such as data leakage often hinder the reproducibility of ML-based research [40]. This is due to the growing number of non-experts employing machine learning across different research fields [24], which is fuelled by the ease of application of ML libraries and no-code off-the-shelf AI tools. In essence, data leakage happens when data, on which the ML model should not be trained, leaks into the training process. Data leakage can be categorized into 3 subcategories [40]:
@@
    Use of non-legitimate data. For example, when the use of antihypertensive drugs is used to predict hypertension. This data is non-legitimate, since it would not be available in a real-world scenario and cannot realistically be used to predict hypertension for a new patient.
@@
    Test set is not drawn from the distribution of scientific interest. There are three possible variants: (i) temporal leakage, which is problematic for ML models that attempt to predict future outcomes, i.e., when some training samples have a later timestamp than samples available in the test set, (ii) the training and test data are not independent of each other, e.g., there should not be samples in the training and test data that are drawn from the same person, and (iii) the test set is not chosen selectively; for instance, if the model is solely evaluated on data, on which it performs well.
@@
Bias in ML refers to the error introduced by approximating a real-world problem, which may be complex, by a simplified model, often leading to systematic deviations from the real world. Furthermore, biases can arise when the model contains imbalances or reflects existing societal biases [47]. ML models, which are subject to bias, are prone to generalization issues, and are therefore potentially problematic for ML reproducibility. There are eight bias variants that can arise during the data handling phase of ML development [67]: (i) selection bias - using data not being representative of the target group, (ii) exclusion bias - excluding particular data samples based on the belief that they are unimportant, (iii) measurement bias - favoring certain measurement results, (iv) recall bias - labeling similar data samples differently, (v) survey bias - introducing data issues stemming from data collection surveys, (vi) confirmation bias - favoring information, which confirms previous beliefs, (vii) prejudice bias - including human-related prejudices in training data, and (viii) algorithmic bias - replicating or amplifying biases by the inner workings of the algorithms. Bias, such as selection bias, often leads to the issue of validity shrinkage in Biomedical science research [38]. For example, in obesity and nutritional research, ML is used to predict obesity, heart rate, or the risk of a heart attack, based on data of an individual. Here, validity shrinkage refers to the issue that a predictive model trained on a subset of data will most probably not perform well on new samples. The difference between predictive performance on known data and new data, however, is most often not accounted for in nutritional science, and therefore also leads to performance claims that cannot be reproduced [38].
@@ 3.4 R4 Experiment
Inherent nondeterminism in ML models means that results can vary between test runs, even with identical code, data, and hyperparameters. This variation occurs from unanticipated influences during training, such as random parameter initialization, stochastic optimization, and random data subsampling (e.g., in k-fold cross-validation) [65, 44]. Neural networks are especially known for their inherent nondeterminism, leading to varied computational outcomes in multiple reruns due to numerous sources of randomness during training [1]. In some cases, inherent non-determinism can cause such large variations that reruns not only yield slightly different outcomes, but also lead to significant fluctuations in the performance of an ML model [1]. This issue is exacerbated when other sources of variation, such as different hyperparameters, are introduced to the ML model. In such cases, the impact can be magnified, and it is often observed that minor changes in hyperparameters can result in significant performance loss [11]. For example, a review of the status quo of reproducibility in NLP research highlights some core issues with non-determinism that are exemplary for ML research [11]. Simply rerunning the original code during a reproduction can lead to different computational outcomes compared to the originally reported ones. Reinforcement learning, a subfield of ML, is particularly susceptible to these reproducibility issues, partially because of additional sources of non-determinism in the learning process, such as the reinforcement learning environment or policy [52].
@@
Various studies have demonstrated that hardware differences, such as different GPUs or CPUs, and compiler settings can lead to different computational outcomes [34]. Additionally, a comparison between the same ML algorithm with fixed random seeds executed using PyTorch [^1] https://pytorch.org/ and TensorFlow [^2] https://www.tensorflow.org/ resulted in different performances [60]. A comparison of the results of an ML experiment between different ML platforms also shows that out-of-the-box reproducibility is not provided. In general, it is questionable to what extent reproducibility can be ensured out-of-the-box [30]. Another important factor is the use of GPUs, which can increase randomness compared to the use of CPUs. This is due to parallel optimization and the use of optimizers in ML frameworks, such as PyTorch and TensorFlow. As a result, some researchers have resorted to solely using CPUs for executing their experiments. However, this comes at the expense of runtime-efficiency [3].
@@
Another barrier to ML reproducibility is the limited access to computational resources, which most recently has become evident in the case of transformer-based large language models (LLMs) [10, 82]. These transformer architectures need a vast amount of data and computational resources, to which most researchers have limited access. Estimates have calculated the costs to reproduce one model to be around \$1 million to \$3.2 million, which largely increases the difficulty of reproducing results [10]. Another study found that the needed computational resources are one of the most significant factors impacting reproducibility [64]. Especially ML models, which require computational clusters for training and optimization, are notably hard to reproduce.